OpenStudy (anonymous):

If it takes Dave 11 hours to close an order and John needs 13 hours how long would it take them together

OpenStudy (anonymous):

13 + 11 ÷ 2

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

\[\frac{11\times 13}{11+13}\]

terenzreignz (terenzreignz):

Really? I'd have thought if they were working together, the answer should be less than if they were working alone

OpenStudy (anonymous):

D: I'm a retard... Okay so @Satellite73 can you explain this to me haha...I should be failing my own math right now...but somehow I am not. o.O

OpenStudy (anonymous):

"close an order" is not really that clear to me, but i am assuming it is one of those problems like "adam takes 13 hours to mow the lawn and eve takes 11 hours. how long would it take for them to paint the room if they work together?"

OpenStudy (anonymous):

@rebeccaskell94 the combined rate is not the average of the two rates

OpenStudy (anonymous):

This I see haha

OpenStudy (anonymous):

dave's rate is \(\frac{1}{11}\) and john's rate is \(\frac{1}{13}\) so their combined rate is \(\frac{1}{11}+\frac{1}{13}=\frac{11+13}{11\times 13}\)

OpenStudy (anonymous):

OHHHH!!!! duh. http://viewfromthedolequeue.files.wordpress.com/2011/04/dunce_cap.jpg <---moi

OpenStudy (anonymous):

then you want how long it takes to complete one job, set \[\frac{11+13}{11\times 13}T=1\] solve via \[t=\frac{11\times 13}{11+13}\]

OpenStudy (anonymous):

of course you still have to compute, i just wrote what you need to compute

OpenStudy (anonymous):

okay! Thank you so much. I shall do my best to remember :D

OpenStudy (anonymous):

you can recreate if for yourself each time, but if you do more than a couple of these you can go right to the answer

OpenStudy (anonymous):

Awesome! Thanks!